A cylindrical tank of radius 3 ft and height 10 ft stands on a flatform 50 ft high. How much work will be done in filling the tank by pumping water from the ground level?


1
Expert's answer
2021-04-29T07:28:08-0400

"\text{Find the volume of the tank:}"

"V = \pi r^2h"

"r = 3ft= 0.9144 m"

"h=10ft = 3.048m"

"V= \pi*0.9144^2*3.048\approx 8m^3"

"\text{Let's calculate the water mass in the tank}"

"m=\rho V= 1000*8 =8000kg"

"\text{Calculate the potential energy of the water in the tank}"

"E_p=mgh_c"

"h_c -\text{height of the center of gravity of the tank }"

"\text{relative to the ground level}"

"h_c= 50+10\/2=55ft=16.764m"

"E_p=mgh_c=8000*9.8*16.764=1314\u00a0297.6J"

"\text{the potential energy of the same mass of water}"

"\text{at ground level will be 0 since the height is 0}"

"E_{p0}=0"

"W =\\Delta P= 1314\u00a0297.6J"

"\text{Answer: } 1314 297.6\ J"